Sam has a bond that has a maturity of 28 years, a 17 year duration, and a yield to maturity of 8.57%. The change in the level of the market interest rate is 0.57%. The modified duration is ________ and the percentage change in price is ________.

28 years; 8.9%

10.41 years; -.57%

13.17 years; 5.7%

15.66 years; -8.9%
Answer #1

Modified Duration=Duration/(1+yield to maturity)=17/1.0857=15.65810076 years

% change=-Modified Duration*change in market rates=-15.65810076*0.57%=-8.925%